Year 0 2 4 Earning & FCF Forecast (Smillions) 1 Sales 2 Growth vs. Prior Year 3 Cost of Goods Sold 4Gross Profit 5 Selling, General & Admin. 6 Depreciation 7 EBIT 8 Less: Income tax at 40% 9 Plus: Depreciation 43300.00% 46800.00% 51600.00% 547 574.3 10.30% 6.00% 5.00% -31360.00% -34570.00%-366.5-384.8 $ 154.40$ 170.30 180.5 189.5 $ (93.60) $ (103.20) -109.4 -114.9 8.10% $(7.00) $(7.50) 596096 62.1 65.2 $ (21.50 $ (23.80 -24.8 -26.1 9 9.5 -770.00000%-1000.0000096-9.9-10.4 8.6-5.6 -4.9 24.6 30.8 33.3 538096 $7.00 7.5 10 Less: Capital Expenditures 11 Less: Increases in NWC 12 Free Cash Flow $ (6.300 25.3Sora Industries has 68 million outstanding​ shares, $ 127 million in​ debt, $ 40 million in​ cash, and the following projected free cash flow for the next four years LOADING...​: a. Suppose​ Sora's revenue and free cash flow are expected to grow at a 4.5 % rate beyond year 4. If​ Sora's weighted average cost of capital is 11.0 %​, what is the value of​ Sora's stock based on this​ information? b.​ Sora's cost of goods sold was assumed to be​ 67% of sales. If its cost of goods sold is actually​ 70% of​ sales, how would the estimate of the​ stock's value​ change? c. ​Let's return to the assumptions of part ​(a​) and suppose Sora can maintain its cost of goods sold at​ 67% of sales.​ However, now suppose Sora reduces its​ selling, general, and administrative expenses from​ 20% of sales to​ 16% of sales. What stock price would you estimate​ now? (Assume no other​ expenses, except​ taxes, are​ affected.) d. ​Sora's net working capital needs were estimated to be​ 18% of sales​ (which is their current level in year​ 0). If Sora can reduce this requirement to​ 12% of sales starting in year​ 1, but all other assumptions remain as in part ​(a​), what stock price do you estimate for​ Sora? ​(Hint​: This change will have the largest impact on​ Sora's free cash flow in year​ 1.)
Answer #1

(a) we have to calculate value of firm using cost of capital as the discount rate, which is 11%.

Value of firm = Present Value of future free cash flows

                     = 25.30/(1.11) + 24.6/(1.11)2 + 30.8/(1.11)3 + 33.3/(1.11)4 + (33.3*1.045/0.11-0.045)/ (1.11)4

                       = 439.87

Value of Equity = 439.87-127(debt value)

                       = 312.87

Value per stock = 312.87/68 = 4.60(approx)

(b) If COGS becomes 70% of sales, All Free cash flows would be reduce by 1.8% of sales ( after tax rate of 3%).

Remaining Solution would be same as in (a)

(c) If Admin expenses reduce by 4% of sales , then all Free cash flows will be increased by 2.4% of sales(after tax rate of 4% )

Year 1 Year 2 Year3 Year 4
Free cash flows 25.3 24.60 30.8 33.30
Add: 2.4% of Sales 11.23 12.38 13.13 13.78
Adjusted FCF 36.53 36.98 43.93 47.08
Present Value 32.91 30 32.12 31.01

value of sora stock would be :

         =   (32.91+30+32.12+31.01+ ((47.08*1.045)/(0.065))/(1.11)4 -127)/68

         = (624.63-127)/68 = 7.318

(d) if working capital needs changes to 12% of sales from 1 year, then

Year 0 Year 1 Year 2 Year3 Year 4
Sales 433 468 516 547 574.30
Op WC-1 -------------------- 77.94 56.16 61.92 65.64
Closing WC -2 77.94(18%*433) 56.16(12%*468) 61.92 65.64 68.92
Change (1-2) ---------------- 21.78(decrease in Working capital) -5.76 -3.72 -3.28
ADD:Free cash flows ------------------ 25.30 24.60 30.80 33.30
ADD: old WC adjustment 6.30 8.60 5.60 4.90
Adjusted Free cash flows 53.38 27.44 32.68 34.92
Present value @ 11% p.a 48.09 22.27 23.89 23

After 4th year, 5th year expected FCF will be 23*1.045=24.035

value of firm at end of 4th year = 24.035/(0.11-0.045) = 369.77

Present value = 369.77/(1.11)4= 243.58

Value of Sora Stock = 48.09+22.27+23.89+23+243.58 -127 = 233.83

Value per Stock = 233.83/68 = 3.44(approx)